OpenStudy (anonymous):

10-16x-x=0 please help! i need to find vertex and asymptote

OpenStudy (lasttccasey):

This is a straight line? Cleans up to be\[y=-17x+10\]

OpenStudy (anonymous):

no like your supposed to get two different x's and then plug them in or something so that you can get the asymptote and the vertex

OpenStudy (lasttccasey):

Straight lines don't have vertexes or asymptotes. Make sure you wrote the equation correctly. There should at least be an x^2 somewhere.

OpenStudy (anonymous):

oh yeah oops its 10-16x-x^2=0

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

well i have to go maybe you can help me later so bye ^-^

OpenStudy (lasttccasey):

Okay cleaning this up into the ax^2+bx+c=0 format we can quickly find the vertex: -x^2-16x+10 Vertex = (h,k) h=-b/2a h=-(-16)/(2*(-1)) h=-8 Plus h back into the equation for the x values to find the k value: -(-8)^2-16(-8)+10=k k=74 Thus the vertex is (-8,74)

OpenStudy (lasttccasey):

And there are no asymptotes for parabolas.